Find 5/63 – (-6/21) Find 5/63 – (-6/21) Find 5/63 – (-6/21)

Find    5/63 – (-6/21)
Find    5/63 – (-6/21)
Find    5/63 – (-6/21)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We can also write -6/21 = -2/7
= 5/63–(-2/7)
We have,
= 5/63 + 2/7
Take the LCM of the denominators of the given rational numbers.
LCM of 63 and 7 is 63
Express each of the given rational numbers with the above LCM as the common denominator.
Now,
(5/63)= [(5×1)/ (63×1)] = (5/63)
(2/7)= [(2×9)/ (7×9)] = (18/63)
Then
= (5/63) + (18/63) ... [∵denominator is same in both the rational numbers]
= (5 + 18)/63
= 23/63

Thanks